OpenStudy (anonymous):

The junior class sold 120 turkey dinner plates and 200 chicken dinner plates for a total of $2,150. The senior class sold 100 turkey plates and 300 chicken plates, raising $2,625. What was the cost of each turkey dinner plate? A$6 B$6.25 C$7 D$7.50 ?????????? help PWEEEEESEEE! (: with almonds&&whip cream&& cranberries &&vanilla icey creamon T O P!!!!!!!!!!!!!!!!!!!!!! D:

OpenStudy (amistre64):

120t + 200c = 2150 ; that is given 100t + 300c = 2625 ; that is given

OpenStudy (amistre64):

how much is a t? ; get rid of the c's and solve really :)

OpenStudy (amistre64):

120t + 200 = 2150 ; *3 100t + 300c = 2625 ; *-2 360t + 600c = 2150(3) = 6450 -200t - 600c = 2625(-2) = -5250 ------------------------------ 160t + 0c = 1200 160t = 1200 t = 1200/160 = 120/16 = 60/8 = 30/4 = 7.50

OpenStudy (anonymous):

thank you SOOOOOOOOOOOOOOO much!!! your the best!!!!!!